# A convergence proof for a finite element discretization of Chorin’s projection method of the incompressible Navier-Stokes equations

### v1

@@ 5.1. Convergence of the numerical scheme
We recall from Remark 5.2 that $\bar{u}_{h}$ is given by the projection of $\widetilde{u}_{h}$ onto $\mathbb{V}_{h}$, specifically $\bar{u}_{h}=\mathcal{P}_{\mathbb{V}_{h}}\widetilde{u}_{h}$. Then, we can estimate the quantity in (5.11) as follows, using that $\mathcal{P}_{\mathbb{V}_{h}}$ is an orthogonal projection:
$\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert_{L^{2}(\Omega)}^{2}dt=\underbrace{\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)-(\bar{u}_{h}(t+\tau)-\bar{u}_{h}(t))\right\rVert_{L^{2}(\Omega)}^{2}dt}_{(i)}\\ +\underbrace{\int_{0}^{T-\tau}\left\lVert\bar{u}_{h}(t+\tau)-\bar{u}_{h}(t)\right\rVert_{L^{2}(\Omega)}^{2}dt}_{(ii)}.$
We let $\epsilon>0$ arbitrary and given. For the first term, we have, using that $\left\lVert\widetilde{u}_{h}-\bar{u}_{h}\right\rVert_{L^{2}}^{2}\leq C\Delta t$ by Lemma 4.2, that there is $h_{1}>0$ (and corresponding $\Delta t$) small enough such that for any $0<h<h_{1}$, we have
$(i)\leq 2\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\bar{u}_{h}(t+\tau)\right\rVert_{L^{2}(\Omega)}^{2}dt+2\int_{0}^{T}\left\lVert\widetilde{u}_{h}(t)-\bar{u}_{h}(t)\right\rVert_{L^{2}(\Omega)}^{2}dt<\frac{\epsilon}{2}.$
For the second term (ii), we use Lemma 5.8 combined with the weak time continuity of $\widetilde{u}_{h}$, Lemma 5.7, the discrete energy inequality, Lemma 5.4 and that $\bar{u}_{h}=\mathcal{P}_{\mathbb{V}_{h}}\widetilde{u}_{h}$ to estimate for any $\eta>0$ and any $0<h<\min\{h_{\eta},h_{1}\}$,
$\begin{split}(ii)&\leq\eta\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert_{L^{2}}^{2}dt+C_{\eta}\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert_{Z}^{2}dt\\ &\leq\eta C_{E_{0},f}+C_{\eta}\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert_{Z}^{2}dt\end{split}$
where $Z:=(H^{1}_{\operatorname{div}}(\Omega)\cap H^{2}(\Omega))^{*}$. Now we first choose $\eta=\epsilon/(4C_{E_{0},f})$ and then since we know from Lemma 5.7 that the second term goes to zero as $\tau\to 0$, we have that for any $0<\tau<\bar{\tau}$ small enough that
$C_{\eta}\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert_{Z}^{2}dt\leq\frac{\epsilon}{4}.$
@@
The result of this lemma, combined with the uniform $L^{2}([0,T];H^{1}_{0}(\Omega))$-bound on the $\widetilde{u}_{h}$ coming from the energy estimate, Lemma 5.4, allow us to apply Simon’s version of the Aubin-Lions-Simon lemma, Theorem B.1, to conclude that a subsequence of $\{\widetilde{u}_{h}\}_{h>0}$ converges strongly in $L^{2}([0,T]\times\Omega)$ to the limit $u$. By Lemma 4.2, this implies that also $\{\bar{u}_{h}\}_{h>0}$ has a strongly convergent subsequence to the same limit $u$. Then the fact that $u$ is a Leray-Hopf solution of the incompressible Navier-Stokes equations follows along the lines of Theorem 4.5 with only slight modifications, or one can also check that Step 2 of Theorem 3.8 in [12] applies.
@@ Appendix A Discrete Grönwall inequality
The following lemma is useful for proving an energy bound in the case that the external source in the fluid equation is nonzero.
@@ Appendix B Aubin-Lions-Simon lemma
We recall Simon’s version of the Aubin-Lions-Simon lemma [27, Theorem 1]:

### v2

@@ 5.1. Convergence of the numerical scheme
We recall from Remark 5.1 that ${u}_{h}$ is given by the projection of $\widetilde{u}_{h}$ onto $\mathbb{V}_{h}$, specifically ${u}_{h}^{m}=\mathcal{P}_{\mathbb{V}_{h}}\widetilde{u}_{h}^{m}$ for all $m\geq 0$ and therefore also that $\mathcal{P}_{\mathbb{V}_{h}}u_{h}=u_{h}$. Then, we can write $\widetilde{u}_{h}=\widetilde{u}_{h}-u_{h}-\mathcal{P}_{\mathbb{V}_{h}}\widetilde{u}_{h}+\mathcal{P}_{\mathbb{V}_{h}}u_{h}+\mathcal{P}_{\mathbb{V}_{h}}\widetilde{u}_{h}$ and estimate the quantity in (5.11) as follows, using triangle inequality and that $\mathcal{P}_{\mathbb{V}_{h}}$ is an orthogonal projection:
$\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert_{L^{2}(\Omega)}^{2}dt$
$\leq 3\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)-(u_{h}(t+\tau)-u_{h}(t))\right\rVert_{L^{2}(\Omega)}^{2}$
$\quad+3\int_{0}^{T-\tau}\left\lVert\mathcal{P}_{\mathbb{V}_{h}}(\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t))-\mathcal{P}_{\mathbb{V}_{h}}(u_{h}(t+\tau)-u_{h}(t))\right\rVert_{L^{2}(\Omega)}^{2}dt$
$\quad+3\int_{0}^{T-\tau}\left\lVert\mathcal{P}_{\mathbb{V}_{h}}(\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t))\right\rVert_{L^{2}(\Omega)}^{2}dt$
$\leq 6\underbrace{\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)-({u}_{h}(t+\tau)-{u}_{h}(t))\right\rVert_{L^{2}(\Omega)}^{2}dt}_{(i)}$
$\quad+3\underbrace{\int_{0}^{T-\tau}\left\lVert\mathcal{P}_{\mathbb{V}_{h}}({\widetilde{u}}_{h}(t+\tau)-{\widetilde{u}}_{h}(t))\right\rVert_{L^{2}(\Omega)}^{2}dt}_{(ii)}.$
We let $\epsilon>0$ arbitrary and given. For the first term, we have, using that $\left\lVert\widetilde{u}_{h}-{u}_{h}\right\rVert_{L^{2}}^{2}\leq C\Delta t$ by (an adaption of) Lemma 4.2, that there is $h_{1}>0$ (and corresponding $\Delta t$) small enough such that for any $0<h<h_{1}$, we have
$(i)\leq 2\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-{u}_{h}(t+\tau)\right\rVert_{L^{2}(\Omega)}^{2}dt+2\int_{0}^{T}\left\lVert\widetilde{u}_{h}(t)-{u}_{h}(t)\right\rVert_{L^{2}(\Omega)}^{2}dt<\frac{\epsilon}{8}.$
For the second term (ii), we use Lemma 4.5 combined with the weak time continuity of $\widetilde{u}_{h}$, Lemma 5.5, the discrete energy inequality, Lemma 5.3 to estimate for any $\eta>0$ and any $0<h<\min\{h_{\eta},h_{1}\}$,
$\begin{split}(ii)&\leq\eta\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert_{H^{1}_{0}}^{2}dt+C_{\eta}\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert_{Z}^{2}dt\\ &\leq\eta C_{E_{0},f}+C_{\eta}\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert_{Z}^{2}dt\end{split}$
where $Z:=(H^{1}_{\operatorname{div}}(\Omega)\cap H^{s}(\Omega))^{*}$. Now we first choose $\eta=\epsilon/(24C_{E_{0},f})$ and then since we know from Lemma 5.5 that the second term goes to zero as $\tau\to 0$, we have that for any $0<\tau<\bar{\tau}$ small enough that
$C_{\eta}\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert_{Z}^{2}dt\leq\frac{\epsilon}{24}.$
@@
The result of this lemma, combined with the uniform $L^{2}([0,T];H^{1}_{0}(\Omega))$-bound on the $\widetilde{u}_{h}$ coming from the energy estimate, Lemma 5.3, allow us to apply Simon’s version of the Aubin-Lions-Simon lemma, Theorem B.1, to conclude that a subsequence of $\{\widetilde{u}_{h}\}_{h>0}$ converges strongly in $L^{2}([0,T]\times\Omega)$ to the limit $u$. By Lemma 4.2, this implies that also $\{{u}_{h}\}_{h>0}$ has a strongly convergent subsequence to the same limit $u$. Then the fact that $u$ is a Leray-Hopf solution of the incompressible Navier-Stokes equations follows along the lines of Theorem 4.7, or one can also check that Step 2 of Theorem 3.8 in [12] applies.
@@ Appendix A Discrete Grönwall inequality
The following standard lemma is useful for proving an energy bound in the case that the external source in the fluid equation is nonzero.
@@ Appendix B Aubin-Lions-Simon lemma
We recall Simon’s version of the Aubin-Lions-Simon lemma [26, Theorem 1]: